20 g of a binary electrolyte (mol.wt. = 100) are dissolved in 500 g of water. The freezing point of the solution is –0.74°C, Kf = 1.86 K.molality–1. The degree of ionisation of the electrolyte is

Grade:11

2 Answers

Arun
6 years ago
Delta T = 1000 *Kf * w/m* 500
0.74 = 1000* 1.86* 20/m*500
on solving
m = 100
Actual molar mass = 100
Hencce degree of ionisation = 0%
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Delta T = 1000 *Kf * w/m* 500
0.74 = 1000* 1.86* 20/m*500
on solving
m = 100
Actual molar mass = 100
Hencce degree of ionisation = 0%

Thanks and Regards